Stansbury Park Observatory Complex 2 (SPOC 2)
Harmons Observatory
Pouring & Finishing the Floor
1st Attempt
24 MAR 2001 through 15 AUG 2001
(click on images for enlarged view)
Images 001-006, Ken Warner's aerial views of the site before the
concrete trucks arrived, 24 MAR 2001.
001

002

003

004

005

006

007
 Ken Harris cutting stakes for forms, 24 MAR 2001. McNeil
photo
008
 Ken Harris surveying for final grade, 24 MAR 2001. McNeil
photo
009
 Bruce Grim & Siegfried Jachmann "head to head" in the footing
trench. Andy Bogdan and Wayne Clarke look on, 24 MAR 2001.
McNeil photo
010
 Workers look for the first concrete truck to arrive, 24 MAR 2001.
McNeil photo
011
 Much anticipated concrete truck #1 arrives, 24 MAR 2001
012
 "1st mud" for the SPOC 2 footings & floor coming down the chute,
24 MAR 2001
013
 First pour for the SPOC 2 footings & floor, 24 MAR 2001.
McNeil photo
014
 President Jachmann supervises the pour, 24 MAR 2001
015
 Ken Harris and Mike Wilson prepare for a pour, 24 MAR 2001.
McNeil photo
016
 Truck #2 begins to pour, 24 MAR 2001
017
 Truck #2 is joined by truck #3, 24 MAR 2001
018
 Yet another pour, 24 MAR 2001. McNeil photo
019
 And another, 24 MAR 2001. McNeil photo
020
 Patrick Wiggins, Roger Butz, Siegfried Jachmann and Wayne Clarke
move the mud, 24 MAR 2001. McNeil photo
021
 Group portrait: Trucks #1, #2 & #3 at the site, 24 MAR 2001
022
 Preparing to scree the floor with the long 2 x 4, 24 MAR 2001
023
 Mike Wilson, John Moreton and Ken Harris work the concrete, 24
MAR 2001. McNeil photo
024
 Crew keeps an eye on Patrick Wiggins (they can't believe he's actually
working!), 24 MAR 2001. McNeil photo
025
 Screeing of the Ealing room floor nearly done, 24 MAR 2001
026
 Andy Bogdan uses a float on the Ealing Room floor while Bruce Grim
finishes screeing the control room floor, 24 MAR 2001
027
 Ealing & control room floors nearly finished, 24 MAR 2001
028
 Screeing the "Big Scope" room floor. Large black & small
white pipes seen sticking up from the floor will carry computer, video
& control cables between the telescopes and the control room, 24 MAR
2001
029
 Andy Bogdan does some touch up work on the control room floor,
24 MAR 2001
030
 Screeing a corner of the "Big Scope" room floor, 24 MAR 2001
031
 Truck #4 makes its contribution, 24 MAR 2001
032
 No man is an island but John Moreton seems to have marooned himself
on a dry patch in the middle of the "Big Scope" room floor, 24 MAR 2001
033
 Ken Harris giving the floor some finishing touches of his own,
24 MAR 2001
034
 Siegfried Jachmann shows how it's done, 24 MAR 2001
035
 Entire floor ready for brushing as storm clouds gather, 24 MAR
2001
036
 Most of the day's crew pose with their work before going to "Advanced
Training", 24 MAR 2001
037
 Workers pose with the newly finished floor, 24 MAR 2001.
McNeil photo
038
 One good looking floor, 24 MAR 2001
039
 Andy Bogdan brushes the floor, 24 MAR 2001
040
 Patrick Wiggins & Scott Crosby leave the newly finished floor,
24 MAR 2001. McNeil photo
041
 Patrick Wiggins keeps with tradition and inscribes the club's initials
and the year, 24 MAR 2001
042
 Patrick Wiggins inscribes "SLAS 2001" in the fresh concrete, 24
MAR 2001. McNeil photo
043
 SLAS's traditional inscription, 24 MAR 2001
044
 Happy campers Bruce Grim and Andy Bogdan pose in front of the results
of their labors, 24 MAR 2001. McNeil photo
045
 Cleaning up the site, 26 MAR 2001
046
 Cleaning up the site, 26 MAR 2001
047
 Patrick Wiggins, Andy Bogdan, Roger Butz, Boyd McNeil and Bruce
Grim agree that no SLAS function is complete without an "Advanced Training
Session", 26 MAR 2001
048
 Wide angle view of the site, 26 MAR 2001
049
 Close up showing the newly installed safety tape and rebar caps,
26 MAR 2001
050
 Bruce Grim, Andy Bogdan & Roger Butz lay electrical cable from
the power pole to the observatory. 14 AUG 2001 McNeil photo
051
 Bruce Grim hands Andy Bogdan the power cable he then ran through
conduit into the observatory. Roger Butz looks on. 14 AUG 2001
McNeil photo
052
 Newly dug trench containing the conduit and cable which will supply
power to the observatory, 15 AUG 2001
|